OpenStudy (anonymous):

can someone make a graph that represents the eqaution y=x^2-4x+4

myininaya (myininaya):

its a parabola so it has shape U or upsidedown U. Since the coefficient of x^2 is positive then it is U. Now find the x intercepts

OpenStudy (anonymous):

ok

myininaya (myininaya):

x^2-4x+4=(x-2)(x-2)=0 implies the x-intercepts are x=2

OpenStudy (anonymous):

also: \[x^2 -4x +4 = (x-2)^2\] might help with finding the x intercept ;)

OpenStudy (anonymous):

can you make a the grpah online and show me it plz

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y%3Dx^2-4x%2B4

OpenStudy (anonymous):

thanks